an aircraft is vertically above a point which is 10 km west and 15 km north of a control tower . if the aircraft is 4000m above the ground ,how far is it from the control tower ?
Answer by Alan3354(69443)

an aircraft is vertically above a point which is 10 km west and 15 km north of a control tower . if the aircraft is 4000m above the ground ,how far is it from the control tower ?
------------
4000m = 4 km
---
d = sqrt(10^2 + 15^2 + 4^2) = sqrt(341) km